# This Version: March 9, 2024

### v1

@@ Appendix B Proofs
The third and final step shows that if the least and the greatest selections of the upper-hemicontinuous fixed point correspondence $\hat{\theta}$ are increasing in $p$, then the fixed points are increasing in $p$. Since under Assumption 4, $\theta$ is a monotonic function of $m,$ therefore, $\mu_{\min}\equiv\theta(\delta_{\inf\mathbb{S}\times\mathbb{X}})$ and $\mu_{\max}\equiv\theta(\delta_{\sup\mathbb{S}\times\mathbb{X}})$, where $\delta_{\mathbb{S}\times\mathbb{X}}$ denotes the degenerate measure on $\mathbb{S}\times\mathbb{X}$ with its mass at $(s,x)$. Therefore, $\mu\succsim\mu_{\text{min }}$ for all $\mu\in\hat{\theta}\left(\mu_{\min}\right)$ and $\mu\precsim\mu_{\max}$ for all $\mu\in\hat{\theta}\left(\mu_{\max}\right)$. Therefore, for every $p\in P$, $\hat{\theta}(\cdot,p):\left[\mu_{\min},\mu_{\max}\right]\rightarrow 2^{\left[\mu_{\min},\mu_{\max}\right]}$. Notice in step 2 that $F(\mu,p)$ is a convex-valued set of fixed points since $T_{\mu,p}$ is convex-valued; the proof of convex-valuedness follows from EP. Therefore, the set of fixed points from $F$ is convex-valued and given $\theta(m)$ is a continuous function, $\hat{\theta}(\mu,p)$ is convex-valued. Hence, from Acemoglu and Jensen (2013), $\hat{\theta}(\cdot,p):\left[\mu_{\min},\mu_{\max}\right]\rightarrow 2^{\left[\mu_{\min},\mu_{\max}\right]}$ is upper hemicontinuous and convex valued and for each fixed value of $\mu\in[\mu_{\min},\mu_{\max}],$ has least and greatest selections and are increasing in $p$ and from Corollary 2 in Milgrom and Roberts (1994) the least and greatest fixed points $\mu^{*}\in\hat{\theta}(\mu^{*},p)$ will be increasing in $p$. Hence, we have proven that the least and greatest inferred models are increasing in the primitives under a positive shock to a regular SMDP. From Theorem 2.8.3 in Topkis (1998), and by treating a change (positive/negative) in $\mu$ as a change in primitive, the remaining part of Theorem 2 follows.
@@
While we have corresponding upper bounds for the utility function, its first derivative, and the difference in the policy functions under the two learning instances, we still need to establish a corresponding upper bound for the total-variation (TV) norm for densities $Q.$ Towards this end, we rely on a lemma by Canonne (2022), that provides an almost stricter bound on the TV-norm between densities. [^55] Our choice of this particular bound over the well-known Pinsker bound is that as the KL divergence gets larger, the Pinsker bound becomes vacuous and exceeds the trivial bound of 1. The BH bound, however, never exceeds 1.

### v2

@@ Appendix B Proofs
The third and final step shows that if the least and the greatest selections of the upper-hemicontinuous fixed point correspondence $\hat{\theta}$ are increasing in $p$, then the fixed points are increasing in $p$. Since under Assumption 4, $\theta$ is a monotonic function of $m,$ therefore, $\mu_{\min}\equiv\theta(\delta_{\inf\mathbb{S}\times\mathbb{X}})$ and $\mu_{\max}\equiv\theta(\delta_{\sup\mathbb{S}\times\mathbb{X}})$, where $\delta_{\mathbb{S}\times\mathbb{X}}$ denotes the degenerate measure on $\mathbb{S}\times\mathbb{X}$ with its mass at $(s,x)$. Therefore, $\mu\succsim\mu_{\text{min }}$ for all $\mu\in\hat{\theta}\left(\mu_{\min}\right)$ and $\mu\precsim\mu_{\max}$ for all $\mu\in\hat{\theta}\left(\mu_{\max}\right)$. Therefore, for every $p\in P$, $\hat{\theta}(\cdot,p):\left[\mu_{\min},\mu_{\max}\right]\rightarrow 2^{\left[\mu_{\min},\mu_{\max}\right]}$. Notice in step 2 that $F(\mu,p)$ is a convex-valued set of fixed points since $T_{\mu,p}$ is convex-valued; the proof of convex-valuedness follows from EP. Therefore, the set of fixed points from $F$ is convex-valued and given $\theta(m)$ is a continuous function, $\hat{\theta}(\mu,p)$ is convex-valued. Hence, from Acemoglu and Jensen (2013), $\hat{\theta}(\cdot,p):\left[\mu_{\min},\mu_{\max}\right]\rightarrow 2^{\left[\mu_{\min},\mu_{\max}\right]}$ is upper hemicontinuous and convex valued and for each fixed value of $\mu\in[\mu_{\min},\mu_{\max}],$ has least and greatest selections and are increasing in $p$ and from Corollary 2 in Milgrom and Roberts (1994) the least and greatest fixed points $\mu^{*}\in\hat{\theta}(\mu^{*},p)$ will be increasing in $p$. Hence, we have proven that the least and greatest inferred models are increasing in the primitives under a positive shock to a regular SMDP. Furthermore, From Theorem 2.8.3 in Topkis (1998), and by treating a change in the beliefs $\mu$ in the usual stochastic order sense as a change in primitive as above, the remaining part of Theorem 2 follows.
@@
While we have corresponding upper bounds for the utility function, its first derivative, and the difference in the policy functions under the two learning instances, we still need to establish a corresponding upper bound for the total-variation (TV) norm for densities $Q.$ Towards this end, we rely on a lemma by Canonne (2022), that provides an almost stricter bound on the TV-norm between densities. [^62] Our choice of this particular bound over the well-known Pinsker bound is that as the KL divergence gets larger, the Pinsker bound becomes vacuous and exceeds the trivial bound of 1. The BH bound, however, never exceeds 1; see Canonne (2022) for further details.